Parrish Inc. Graded HW

At December 31, 2001, Parrish Inc. reported this information on its balance sheet:

Accounts receivable $2,000,000

Less: Allowance for doubtful accounts 120,000

During 2002, the company had the following transactions:

1. Sales on account $5,200,000

2. Sales returns and allowances 80,000

3. Collections of accounts receivables 4,600,000

4. Write-offs of accounts receivable deemed uncollectible 130,000

5. Recovery of bad debts previously written off 50,000 Required:

a. Prepare the journal entries to record each of these five transactions.

b. Prepare the journal entry to record the bad debt expense assuming that the company assumes its bad debts at 2% of net sales.

c. Enter the January 1, 2002, balances in Accounts Receivable and Allowance for doubtful accounts, post the entries to the two accounts (T accounts) and determine the balances.

d. Show the impact of these transactions on the 2002 income statement and balance sheet.

e. Compute the receivables turnover ratio and the average collection period for the company for 2002. The average collection period is 75 days in the industry. If you were a banker would you extend a loan to the company, why or why not.
Answer #1

a.

No Account Titles and Explanation Debit Credit
1 Accounts receivable 5,200,000
Sales revenue 5,200,000
2 Sales returns and allowances 80,000
Accounts receivable 80,000
3 Cash 4,600,000
Accounts receivable 4,600,000
4 Allowance for doubtful accounts 130,000
Accounts receivable 130,000
5 Accounts receivable 50,000
Allowance for doubtful accounts 50,000
Cash 50,000
Accounts receivable 50,000

b.

Account Titles and Explanation Debit Credit
Bad debt expense 102400
Allowance for doubtful accounts 102400
[5,200,000-80,000] *2%

c.

Accounts Receivable
Beg Bal 2,000,000 2 80000
1 5,200,000 3 4,600,000
5 50,000 4 130,000
5 50,000
End bal 2,390,000
Allowance for doubtful accounts
4 130,000 Beg Bal 120,000
5 50,000
End Bal 40,000

d.

Income Statement
Sales revenue 5,200,000
Expenses
Bad debt expenses 130,000
Balance Sheet
Current Assets
Accounts receivable 2,390,000
Less: Allowance for doubtful accounts 40,000

e. Receivables turn over ratio= Net credit sales / Avg receivables = ( 5,200,000- 80,000) / [ (2,390,000+ 2,000,000)/2] = 2.33 times

average collection period = 365 / Receivables turn over ratio = 365/ 2.33 = 156

No, I would not extend the loan if I were a banker since the average collection period is way higher than the industry standards.